# A complete classification of the zero-dimensional homogeneous spaces under determinacy

### v2

@@ 5.2. At most one class in every pair has the separation property
Theorem 5.6 (Van Wesep).
Proof.
Since $\omega^{\omega}\times\omega^{\omega}\approx\omega^{\omega}$, using Lemma 5.4 we can fix $(A_{0},A_{1})\in\mathbf{\Gamma}\times\mathbf{\Gamma}$ such that for every $(B_{0},B_{1})\in\mathbf{\Gamma}\times\mathbf{\Gamma}$ there exists a continuous $f:\omega^{\omega}\longrightarrow\omega^{\omega}$ such that $f^{-1}[A_{i}]=B_{i}$ for each $i$. Define the game $\mathsf{G}_{0}$ by declaring that Player II wins if the following conditions are satisfied:
  • $x\in A_{0}\setminus A_{1}\rightarrow y\in A_{1}\setminus A_{0}$,
  • $x\in A_{1}\setminus A_{0}\rightarrow y\in A_{0}\setminus A_{1}$,
  • $y\notin A_{0}\cap A_{1}$,
where $x\in\omega^{\omega}$ and $y\in\omega^{\omega}$ denote the sequences played by Player I and Player II respectively. Similarly, define the game $\mathsf{G}_{1}$ by declaring that Player II wins if the following conditions are satisfied:
  • $x\in A_{0}\setminus A_{1}\rightarrow y\in A_{0}\setminus A_{1}$,
  • $x\in A_{1}\setminus A_{0}\rightarrow y\in A_{1}\setminus A_{0}$,
  • $y\in A_{0}\cup A_{1}$.
Claim 1. If Player II has a winning strategy in $\mathsf{G}_{1}$ then $\mathbf{\Gamma}$ has the second separation property.
Proof of Claim 1. Let $\tau$ be a winning strategy for Player II in $\mathsf{G}_{0}$. Set $g=f_{\tau}$, and observe that $g$ is continuous by Proposition 2.1. Define $A_{i}^{\prime}=g^{-1}[\omega^{\omega}\setminus A_{1-i}]$ for $i\in 2$, and observe that each $A_{i}^{\prime}\in\widecheck{\mathbf{\Gamma}}$. Furthermore, since $\tau$ is winning, one sees that $A_{0}\setminus A_{1}\subseteq A_{0}^{\prime}$, $A_{1}\setminus A_{0}\subseteq A_{1}^{\prime}$, and $A_{0}^{\prime}\cap A_{1}^{\prime}=\varnothing$. The property of $(A_{0},A_{1})$ stated at the beginning of the proof then yields the desired result. $\blacksquare$
Claim 2. If Player II has a winning strategy in $\mathsf{G}_{0}$ then $\widecheck{\mathbf{\Gamma}}$ has the second separation property.
Proof of Claim 2. This is similar to the proof of Claim 1. $\blacksquare$
Given Claims 1 and 2, and the assumption that $\mathsf{Det}(\mathbf{\Sigma}(\omega^{\omega}))$ holds, it remains to show that Player I cannot have winning strategies for both $\mathsf{G}_{0}$ and $\mathsf{G}_{1}$. Assume, in order to get a contradiction, that this is not the case. Fix a winning strategy $\sigma_{i}$ for Player I in $\mathsf{G}_{i}$ for $i\in 2$. Set $f_{i}=f_{\sigma_{i}}$ for $i\in 2$, and observe that each $f_{i}$ is continuous by Proposition 2.1.
It is clear from the definition of $\mathsf{G}_{0}$ that for every $y\in\omega^{\omega}$ at least one of the following conditions will hold:
  • $f_{0}(y)\in A_{0}\setminus A_{1}$ and $y\notin A_{1}\setminus A_{0}$,
  • $f_{0}(y)\in A_{1}\setminus A_{0}$ and $y\notin A_{0}\setminus A_{1}$,
  • $y\in A_{0}\cap A_{1}$.
Therefore, the following conditions will hold for every $y\in\omega^{\omega}$:
  • $y\in A_{0}\setminus A_{1}\rightarrow f_{0}(y)\in A_{0}\setminus A_{1}$,
  • $y\in A_{1}\setminus A_{0}\rightarrow f_{0}(y)\in A_{1}\setminus A_{0}$,
  • $y\in\omega^{\omega}\setminus(A_{0}\cup A_{1})\rightarrow f_{0}(y)\in(A_{0}\setminus A_{1})\cup(A_{1}\setminus A_{0})$.
By a similar argument, the following conditions will hold for every $y\in\omega^{\omega}$:
  • $y\in A_{0}\setminus A_{1}\rightarrow f_{1}(y)\in A_{1}\setminus A_{0}$,
  • $y\in A_{1}\setminus A_{0}\rightarrow f_{1}(y)\in A_{0}\setminus A_{1}$,
  • $y\in A_{0}\cap A_{1}\rightarrow f_{1}(y)\in(A_{0}\setminus A_{1})\cup(A_{1}\setminus A_{0})$.
Set $z^{\ast}_{n}=\lambda(n,z)$ for $(n,z)\in\omega\times 2^{\omega}$, where $\lambda$ is the function obtained by applying Lemma 5.3 with $\Omega=2$ to $\sigma_{0}$ and $\sigma_{1}$. Given $S\subseteq\omega^{\omega}$, set
$S^{\ast}=\{z\in 2^{\omega}:z^{\ast}_{0}\in S\}.$
Notice that $S^{\ast}\subseteq\mathbf{\Sigma}(2^{\omega})$ whenever $S\subseteq\mathbf{\Sigma}(\omega^{\omega})$, since the function $\lambda(0,-)$ is continuous. In particular, all sets of the form $S^{\ast}$ considered in the remainder of this proof will have the Baire property by Corollary 2.7. Also observe that $(S\cup T)^{\ast}=S^{\ast}\cup T^{\ast}$ for all $S,T\subseteq 2^{\omega}$.
Make the following definitions:
  • $C=(A_{0}\setminus A_{1})\cup(A_{1}\setminus A_{0})$,
  • $M=(A_{0}\cap A_{1})\cup(\omega^{\omega}\setminus(A_{0}\cup A_{1}))$.
Claim 3. $M^{\ast}$ is meager.
Proof of Claim 3. Assume that $M^{\ast}$ is non-meager. By [Ke, Proposition 8.26], we can fix $s\in 2^{<\omega}$ such that $\{z\in 2^{\omega}:(s^{\frown}z)^{\ast}_{0}\in M\}$ is comeager. Using Lemma 5.3 and the properties of the functions $f_{i}$ listed above, it is easy to realize that $z^{\ast}_{0}\in C$ implies $(i^{\frown}z)^{\ast}_{0}\in C$ for each $i$. It follows that $(s^{\frown}z)^{\ast}_{0}\in M$ implies $z^{\ast}_{0}\in M$, hence $M^{\ast}\supseteq\{z\in 2^{\omega}:(s^{\frown}z)^{\ast}_{0}\in M\}$ is comeager. In particular, either $(A_{0}\cap A_{1})^{\ast}$ or $(\omega^{\omega}\setminus(A_{0}\cup A_{1}))^{\ast}$ is non-meager. We will assume that $(A_{0}\cap A_{1})^{\ast}$ is non-meager, since the other case is similar. Notice that if $z^{\ast}_{0}\in A_{0}\cap A_{1}$ then $(1^{\frown}z)^{\ast}_{0}\in C$. This shows that $\{z\in 2^{\omega}:(1^{\frown}z)^{\ast}_{0}\in C\}\supseteq(A_{0}\cap A_{1})^{\ast}$ is non-meager, hence $C^{\ast}$ is non-meager. This contradicts the fact that $M^{\ast}$ is comeager. $\blacksquare$
By Claim 3, either $(A_{0}\setminus A_{1})^{\ast}$ or $(A_{1}\setminus A_{0})^{\ast}$ is non-meager. Assume that $(A_{0}\setminus A_{1})^{\ast}$ is non-meager, since the other case is symmetrical. By [Ke, Proposition 8.26], we can fix $s\in 2^{<\omega}$ such that $\{z\in 2^{\omega}:(s^{\frown}z)^{\ast}_{0}\in A_{0}\setminus A_{1}\}$ is comeager. First assume that $s$ contains an odd number of ones. Then $z^{\ast}_{0}\in A_{0}\setminus A_{1}$ implies $(s^{\frown}z)^{\ast}_{0}\notin A_{0}\setminus A_{1}$. This means that $(A_{0}\setminus A_{1})^{\ast}\subseteq 2^{\omega}\setminus\{z\in 2^{\omega}:(s^{\frown}z)^{\ast}_{0}\in A_{0}\setminus A_{1}\}$ is meager, which is a contradiction.
To conclude the proof, assume that $s$ contains an even number of ones. Then $(A_{1}\setminus A_{0})^{\ast}\subseteq 2^{\omega}\setminus\{z\in 2^{\omega}:(s^{\frown}z)^{\ast}_{0}\in A_{0}\setminus A_{1}\}$ is meager. It follows from Claim 3 that $(A_{0}\setminus A_{1})^{\ast}$ is comeager. Since $\{z\in 2^{\omega}:(1^{\frown}z)^{\ast}_{0}\in A_{1}\setminus A_{0}\}\supseteq(A_{0}\setminus A_{1})^{\ast}$, this shows that $(A_{1}\setminus A_{0})^{\ast}$ is non-meager, which is a contradiction. ∎
5.3. At least one class in every pair has the separation property
The following is the main result of [St2]. Notice that, unlike in the previous section, the second separation property will be entirely bypassed.
Theorem 5.7 (Steel).
Proof.
Call $(E_{0},E_{1})$ an inseparable pair if the following conditions hold:
  • Either $E_{0},E_{1}\in\mathbf{\Gamma}$ or $E_{0},E_{1}\in\widecheck{\mathbf{\Gamma}}$,
  • $E_{0}\cap E_{1}=\varnothing$,
  • There exists no $C\in\Delta(\mathbf{\Gamma})$ such that $E_{0}\subseteq C\subseteq\omega^{\omega}\setminus E_{1}$.
Given $D_{0},D_{1},E_{0},E_{1}\subseteq\omega^{\omega}$, define the game $\mathsf{G}(D_{0},D_{1},E_{0},E_{1})$ by declaring that Player I wins if
$y\in D_{i}\rightarrow x\in E_{i}\text{ for every }i\in 2.$
where $x\in\omega^{\omega}$ and $y\in\omega^{\omega}$ denote the sequences played by Player I and Player II respectively.
Claim. Let $D_{0},D_{1}\subseteq\omega^{\omega}$, and let $(E_{0},E_{1})$ be an inseparable pair. Assume that the following conditions hold:
  • Either $D_{0},D_{1}\in\mathbf{\Gamma}$ or $D_{0},D_{1}\in\widecheck{\mathbf{\Gamma}}$,
  • $D_{0}\cap D_{1}=\varnothing$.
Then Player I has a winning strategy in $\mathsf{G}(D_{0},D_{1},E_{0},E_{1})$.
Proof of the claim. Assume, in order to get a contradiction, that Player I has no winning strategy. By the assumption that $\mathsf{Det}(\mathbf{\Sigma}(\omega^{\omega}))$ holds, we can fix a winning strategy $\tau$ for Player II. Set $g=f_{\tau}$, and observe that $g$ is continuous by Proposition 2.1. Using the fact that $\tau$ is winning, one sees that $g[E_{i}]\subseteq D_{1-i}$ for each $i$ and $g[\omega^{\omega}]\subseteq D_{0}\cup D_{1}$. It follows that $E_{0}\subseteq g^{-1}[D_{1}]\subseteq\omega^{\omega}\setminus E_{1}$ and that $g^{-1}[D_{1}]=g^{-1}[\omega^{\omega}\setminus D_{0}]\in\Delta(\mathbf{\Gamma})$, which is a contradiction. $\blacksquare$
Assume, in order to get a contradiction, that $(A_{0},A_{1})\in\mathbf{\Gamma}\times\mathbf{\Gamma}$ and $(B_{0},B_{1})\in\widecheck{\mathbf{\Gamma}}\times\widecheck{\mathbf{\Gamma}}$ are both inseparable pairs. By the claim, we can fix a winning strategy $\sigma$ for Player I in $\mathsf{G}(A_{0},A_{1},B_{0},B_{1})$. Set $f=f_{\sigma}$, and observe that $f$ is continuous by Proposition 2.1. Define $C_{i}=f^{-1}[B_{i}]$ for $i\in 2$. Notice that $f[A_{i}]\subseteq B_{i}$ for each $i$ because $\sigma$ is winning, hence $A_{i}\subseteq C_{i}$ for each $i$. Furthermore, it is clear that $C_{i}\in\widecheck{\mathbf{\Gamma}}$ for each $i$ and $C_{0}\cap C_{1}=\varnothing$.
By applying the claim three more times, we can fix winning strategies $\sigma_{0}$, $\sigma_{1}$ and $\sigma_{2}$ for Player I in the games $\mathsf{G}(A_{1},A_{0},A_{0},A_{1})$, $\mathsf{G}(A_{0},\omega^{\omega}\setminus C_{0},A_{0},A_{1})$ and $\mathsf{G}(\omega^{\omega}\setminus C_{1},A_{1},A_{0},A_{1})$ respectively. Set $f_{i}=f_{\sigma_{i}}$ for $i\in 3$, and observe that each $f_{i}$ is continuous by Proposition 2.1. Furthermore, it is easy to realize that the following inclusions hold:
  • $f_{0}[A_{1}]\subseteq A_{0}$ and $f_{0}[A_{0}]\subseteq A_{1}$,
  • $f_{1}[A_{0}]\subseteq A_{0}$ and $f_{1}[\omega^{\omega}\setminus C_{0}]\subseteq A_{1}$,
  • $f_{2}[\omega^{\omega}\setminus C_{1}]\subseteq A_{0}$ and $f_{2}[A_{1}]\subseteq A_{1}$.
Set $z^{\ast}_{n}=\lambda(n,z)$ for $(n,z)\in\omega\times 3^{\omega}$, where $\lambda$ is the function obtained by applying Lemma 5.3 with $\Omega=3$ to $\sigma_{0}$, $\sigma_{1}$ and $\sigma_{2}$. Given $S\subseteq 3^{\omega}$, set
$S^{\ast}=\{z\in 3^{\omega}:z^{\ast}_{0}\in S\}.$
Notice that $S^{\ast}\subseteq\mathbf{\Sigma}(3^{\omega})$ whenever $S\subseteq\mathbf{\Sigma}(\omega^{\omega})$, since the function $\lambda(0,-)$ is continuous. In particular, all sets of the form $S^{\ast}$ considered in the remainder of this proof will have the Baire property by Corollary 2.7. Also observe that $(S\cup T)^{\ast}=S^{\ast}\cup T^{\ast}$ for all $S,T\subseteq 2^{\omega}$.
First assume that $(A_{0}\cup A_{1})^{\ast}$ is comeager. This means that we can fix $i\in 2$ such that $A_{i}^{\ast}$ is non-meager. By [Ke, Proposition 8.26], we can fix $s\in 3^{<\omega}$ such that $\{z\in 3^{\omega}:(s^{\frown}z)^{\ast}_{0}\in A_{i}\}$ is comeager. Using Lemma 5.3 and the properties of the functions $f_{i}$ listed above, one sees that if $s$ contains an odd number of zeros, then $z^{\ast}_{0}\in A_{i}$ implies $(s^{\frown}z)^{\ast}_{0}\in A_{1-i}$. On the other hand, if $s$ contains an even number of zeros, then $z^{\ast}_{0}\in A_{i}$ implies $(s^{\frown}0^{\frown}z)^{\ast}_{0}\in A_{1-i}$. In either case, it is easy to realize that $\{z\in 3^{\omega}:(s^{\frown}z)^{\ast}_{0}\in A_{1-i}\}$ is non-meager, which is a contradiction.
Finally, assume that $(A_{0}\cup A_{1})^{\ast}$ is non-comeager. By [Ke, Proposition 8.26], we can fix $s\in 3^{<\omega}$ such that $\{z\in 3^{\omega}:(s^{\frown}z)^{\ast}_{0}\notin A_{0}\cup A_{1}\}$ is comeager. Since $(s^{\frown}z)^{\ast}_{0}\notin A_{0}\cup A_{1}$ implies $z^{\ast}_{0}\notin A_{0}\cup A_{1}$, it follows that $\{z\in 3^{\omega}:z^{\ast}_{0}\notin A_{0}\cup A_{1}\}$ is comeager. Since $(\omega^{\omega}\setminus C_{0})\cup(\omega^{\omega}\setminus C_{1})=\omega^{\omega}$, we can fix $i\in 2$ such that $(\omega^{\omega}\setminus C_{i})^{\ast}$ is non-meager. Since $z^{\ast}_{0}\notin C_{i}$ implies $((i+1)^{\frown}z)^{\ast}_{0}\in A_{0}\cup A_{1}$, one sees that $(A_{0}\cup A_{1})^{\ast}$ is non-meager. This contradiction concludes the proof. ∎
5.4. The solution to the “pair of socks” problem
We are finally in a position to obtain the generalization promised in Section 5.1.
@@ 5.4. The solution to the “pair of socks” problem
As a first application, we will show that expansions preserve the separation property. This result will be useful in the proof of Corollary 9.14.
@@
Let $\mathbf{\Gamma}(Z)\in\mathsf{NSD}_{\mathbf{\Sigma}}(Z)$, where $\mathbf{\Gamma}\in\mathsf{NSD}_{\mathbf{\Sigma}}(\omega^{\omega})$, and assume that $\mathbf{\Gamma}(Z)$ has the separation property. Pick disjoint $A,B\in\mathbf{\Gamma}(Z)^{(\xi)}$. By Lemma 3.31, we can fix a zero-dimensional Polish space $W$ and a $\mathbf{\Sigma}^{0}_{1+\xi}$-measurable bijection $f:Z\longrightarrow W$ such that $f[A],f[B]\in\mathbf{\Gamma}(W)$. Observe that $\mathbf{\Gamma}$ must have the separation property, otherwise Lemma 5.2 and Theorem 5.8 would easily yield a contradiction. Therefore $\mathbf{\Gamma}(W)$ has the separation property by Lemma 5.2. This means that we can find $C\in\Delta(\mathbf{\Gamma}(W))$ such that $f[A]\subseteq C\subseteq W\setminus f[B]$. Notice that $f^{-1}[C]\in\Delta\left(\mathbf{\Gamma}(Z)^{(\xi)}\right)$ by Lemmas 3.29 and 3.30. Since clearly $A\subseteq f^{-1}[C]\subseteq Z\setminus B$, this concludes the proof. ∎
Chapter 6 Wadge theory: type
The purpose of this chapter is to introduce the notions of type and separated unions, discuss their basic properties, and characterize the classes of small type. Loosely speaking, these are the classes that can be obtained by applying the operation of separated union to a countable collection of classes of lower complexity (see Theorem 6.11). Much like the separation property, type and separated unions are not relevant to the uniqueness theorem, but they will be important for the existence results, because of the role that they play in obtaining the closure properties of Section 7.5.
@@ 6.1. Definition and basic properties
The following notion is essentially due to Louveau (see [Lo1, Definition 1.7] and the remark that precedes it), although his “official” definition is of a (rather complicated) recursive nature, and it is limited to the Borel context. This is one of the fundamental notions needed to state our classification results.111 It would have been more elegant to simply define $t(\mathbf{\Gamma})$ as the cofinality of $\mathbf{\Gamma}$ in $\mathsf{NSD}^{(\xi)}(Z)$. However, we preferred to remain consistent with the original definition.
@@
$\left.\begin{array}[]{lcl}&&t(\mathbf{\Gamma})=\left\{\begin{array}[]{ll}0&\text{if }\mathbf{\Gamma}\text{ is minimal in $\mathsf{NSD}^{(\xi)}(Z)$,}\\ 1&\text{if }\mathbf{\Gamma}\text{ is a successor in $\mathsf{NSD}^{(\xi)}(Z)$,}\\ 2&\text{if }\mathbf{\Gamma}\text{ is a countable limit in $\mathsf{NSD}^{(\xi)}(Z)$,}\\ 3&\text{otherwise.}\\ \end{array}\right.\end{array}\right.$
@@
    $\left.\begin{array}[]{lcl}&&t(\mathbf{\Gamma})=\left\{\begin{array}[]{ll}1&\text{if }\mathbf{\Gamma}\text{ is a successor in $\mathsf{NSD}(Z)$,}\\ 2&\text{if }\mathbf{\Gamma}\text{ is a countable limit in $\mathsf{NSD}(Z)$,}\\ 3&\text{otherwise.}\\ \end{array}\right.\end{array}\right.$
We conclude this section by showing that expansions leave the type unchanged (see Corollary 6.4 for the precise statement).
@@ 6.2. Separated unions
The following notion will ultimately allow us to explicitly describe the non-selfdual Wadge classes that are successors or countable limits (see Section 6.3).222 According to Wadge, this notion had already been identified by “the classical descriptive set theorists” (see [Wa2, page 187]). For some insight on the relationship between separated unions and separated differences, see Lemma 7.16.
@@
Then $\mathbf{\Gamma}\in\mathsf{NSD}(Z)$.333 However, it is not necessarily true that $\mathbf{\Gamma}\subseteq\mathbf{\Sigma}(Z)$.
@@
Recall that, by Theorem 5.8, exactly one of the Wadge classes in a non-selfdual pair has the separation property. The following result tells us which one, at least in certain cases.

### v3

@@ 3.8. The separation property and the “pair of socks” problem
We are finally in a position to completely solve the “pair of socks” problem.
@@
As a first application, we will show that expansions preserve the separation property. This result will be useful in the proof of Corollary 6.13. Other applications of Theorem 3.50 will be given in the proofs of Corollaries 6.14 and 6.16.
@@
Let $\mathbf{\Gamma}(Z)\in\mathsf{NSD}_{\mathbf{\Sigma}}(Z)$, where $\mathbf{\Gamma}\in\mathsf{NSD}_{\mathbf{\Sigma}}(\omega^{\omega})$, and assume that $\mathbf{\Gamma}(Z)$ has the separation property. Pick disjoint $A,B\in\mathbf{\Gamma}(Z)^{(\xi)}$. By Lemma 3.27, we can fix a zero-dimensional Polish space $W$ and a $\mathbf{\Sigma}^{0}_{1+\xi}$-measurable bijection $f:Z\longrightarrow W$ such that $f[A],f[B]\in\mathbf{\Gamma}(W)$. Observe that $\mathbf{\Gamma}$ must have the separation property, otherwise Lemma 3.45 and Theorem 3.50 would easily yield a contradiction. Therefore $\mathbf{\Gamma}(W)$ has the separation property by Lemma 3.45. This means that we can find $C\in\Delta(\mathbf{\Gamma}(W))$ such that $f[A]\subseteq C\subseteq W\setminus f[B]$. Notice that $f^{-1}[C]\in\Delta\left(\mathbf{\Gamma}(Z)^{(\xi)}\right)$ by Lemmas 3.25 and 3.26. Since clearly $A\subseteq f^{-1}[C]\subseteq Z\setminus B$, this concludes the proof. ∎
3.9. Separated differences
Definition 3.52 (Louveau).
Let $Z$ be a space, let $1\leq\eta<\omega_{1}$, let $U_{\mu,n},A_{\mu,n}\subseteq Z$ for $\mu<\eta$ and $n\in\omega$, and let $A^{\ast}\subseteq Z$. Define
(3) $\mathsf{SD}_{\eta}((U_{\mu,n}:\mu<\eta,n\in\omega),(A_{\mu,n}:\mu<\eta,n\in\omega),A^{\ast})=\\ =\bigcup_{\begin{subarray}{c}\mu<\eta\\ n\in\omega\end{subarray}}\left(A_{\mu,n}\cap U_{\mu,n}\setminus\bigcup_{\begin{subarray}{c}\mu^{\prime}<\mu\\ m\in\omega\end{subarray}}U_{\mu^{\prime},m}\right)\cup\left(A^{\ast}\setminus\bigcup_{\begin{subarray}{c}\mu<\eta\\ n\in\omega\end{subarray}}U_{\mu,n}\right).$
Lemma 3.53.
Lemma 3.54.
Separated differences are important because they make it possible to give concrete descriptions of the non-selfdual Wadge classes of level 0. This is made precise by the following result, which is essentially due to Louveau (it can be easily derived using [CMM2, Proposition 19.4 and Theorem 22.2]). See also the first paragraph of §4.2 to clarify the meaning of $\mathbf{\Delta}$.
Theorem 3.55.
  • $\mathbf{\Gamma}\in\mathsf{NSD}(Z)$ and $\ell(\mathbf{\Gamma})=0$,
4. Wadge theory: type
@@ 4.1. Definition and basic properties
The following notion is essentially due to Louveau (see [Lo1, Definition 1.7] and the remark that precedes it), although his “official” definition is of a (rather complicated) recursive nature, and it is limited to the Borel context. This is one of the fundamental notions needed to state our classification results.141414 It would have been more elegant to simply define $t(\mathbf{\Gamma})$ as the cofinality of $\mathbf{\Gamma}$ in $\mathsf{NSD}^{(\xi)}(Z)$. However, we preferred to remain consistent with the original definition.
@@
[…]
@@
[…]
We conclude this subsection by showing that expansions leave the type unchanged (see Corollary 4.4 for the precise statement).
@@ 4.2. Separated unions
The following notion will ultimately allow us to explicitly describe the non-selfdual Wadge classes that are successors or countable limits (see §4.3).151515 According to Wadge, this notion had already been identified by “the classical descriptive set theorists” (see [Wa2, page 187]). For some insight on the relationship between separated unions and separated differences, see Lemma 5.16.
@@
Then $\mathbf{\Gamma}\in\mathsf{NSD}(Z)$.161616 However, it is not necessarily true that $\mathbf{\Gamma}\subseteq\mathbf{\Sigma}(Z)$.
@@
Recall that, by Theorem 3.50, exactly one Wadge classes in every non-selfdual pair has the separation property. The following result tells us which one, at least in certain cases.